# VR-Based Control of Multi-Copter Operation

### v1

@@
Abstract
We aim to use virtual reality (VR) to improve the spatial awareness of pilots by real-time scanning of the environment around the drone using onboard sensors, live streaming of this environment to a VR headset, and rendering a virtual representation of the drone and its environment for the pilot. This way, the pilot can see the immediate environment of the drone up close from a third-person perspective, as opposed to the first-person perspective that most drone cameras provide. This provides much more information about the drone’s surroundings for the pilot while operating the drone than existing teleoperation solutions. Previous solutions using VR have relied upon pre-made designs of the environment, which makes it difficult to adapt to changing environments. Our solution, in contrast, scans the environment as you fly, making it much more flexible for use in unknown environments.
Index Terms:
Human spatial awareness enhancement, human robot interaction, real-time scanning, drone 3D reconstruction, livestreaming.
@@ III-A Hexacopter Model
To localize the hexacopter motion, we define a global coordinate system that is fixed on the ground with orthonormal base vectors $\hat{\mathbf{e}}_{1}$, $\hat{\mathbf{e}}_{2}$, and $\hat{\mathbf{e}}_{3}$. We also define a local (body) coordinate system that is fixed on the hexacopter body, with orthonormal base vectors $\hat{\mathbf{c}}_{1}$, $\hat{\mathbf{c}}_{2}$, and $\hat{\mathbf{c}}_{3}$. We can use the $3-2-1$ Euler angle standard to characterize the orientation of the hexacopter at any time, where we use $\phi$, $\theta$, and $\psi$ to denote the hexacopter’s roll, pitch, and yaw angles, respectively. The angular velocity of the hexacopter is expressed with respect to the body coordinate system by
@@ III-B Controller
$\hat{\mathbf{c}}_{1,des}=\frac{\left(\hat{\mathbf{e}}_{3}\times\hat{\mathbf{h}}\right)\times\hat{\mathbf{c}}_{3,des}}{\left|\left|\left(\hat{\mathbf{e}}_{3}\times\hat{\mathbf{h}}\right)\times\hat{\mathbf{c}}_{3,des}\right|\right|},$ (6b)
@@
$\hat{\mathbf{h}}=\frac{(\hat{\mathbf{c}}_{1}\cdot\hat{\mathbf{e}}_{1})\hat{\mathbf{e}}_{1}+(\hat{\mathbf{c}}_{1}\cdot\hat{\mathbf{e}}_{2})\hat{\mathbf{e}}_{2}}{||(\hat{\mathbf{c}}_{1}\cdot\hat{\mathbf{e}}_{1})\hat{\mathbf{e}}_{1}+(\hat{\mathbf{c}}_{1}\cdot\hat{\mathbf{e}}_{2})\hat{\mathbf{e}}_{2}||}.$ (7)
@@ IV-B Software Setup
To measure the latency of multi-copter tracking information, we used timestamped ROS2 messages, which are sent to the headset and echoed back to the VR server. We can then divide the round-trip latency by two to get the one-way latency. Rendering latency can be considered negligible, so this value can be considered an end-to-end latency measurement.
@@
Figure 5: Comparison of latency with commercial Offering
Figure 6: Graph of tracking latency. The highlighted portion is the period where mapping was enabled.
Figure 7: Screenshot of partially reconstructed environment, hexacopter, and video stream
@@
    Portability/Flexibility - Our WebXR-based approach will work on any VR headset that supports WebXR now and in the future.
@@
    Bandwidth Utilization - Our compressed video streaming and binary mesh format use up as little bandwidth as possible while quickly transferring all necessary.
@@
    Security - All data transferred between front and back ends is encrypted with TLS.
IV-C Limitations
@@ V Experimental Results
Our test pilot, while only interacting with the system, was able to take off, fly to a wall, stop and land safely, without having a direct visual line of sight to the hexacopter. They were able to recognize the position of the wall by seeing the virtual representation of the wall that was scanned during the flight.
Furthermore, the average latency of the hexacopter tracking data during the duration of the test was 41.9 ms, with a median of 28.7 ms. This means that in the virtual environment, the virtual drone’s position and orientation are being accurately represented in real time, providing visual feedback to pilots equivalent to what they would receive through direct line of sight.
The system clearly demonstrated that the environment can be mapped and displayed to the pilot in real time, as shown in Fig. 7 and the accompanying video. The measured latency between the system sending the data and it being displayed in the headset (see Fig. 5) had a median below commercial FPV drone offerings [26]. Our system’s average latency is significantly higher than the median due to significant latency spikes that can occur when mapping a large environment, as shown in the highlighted region in Fig. 6.

### v2

@@
Abstract—We present a VR-based teleoperation system for multirotor flight that renders a third-person view (TPV) of the vehicle together with a live 3D reconstruction of its surroundings. The system runs on an embedded GPU (Jetson Orin NX) with ROS2/WebXR integration and streams geometry and video to a headset for closed-loop control in previously unmapped spaces. We implement a first-person video (FPV) baseline and perform matched trials with two pilots in unmapped indoor spaces. Quantitative metrics are reported from repeated trials with one pilot (N=8N{=}8). TPV achieved task time comparable to FPV while improving proximal obstacle awareness (minimum obstacle distance $0.20\,\mathrm{m}$) and reducing contacts. These results indicate that TPV can preserve control quality while exposing hazards less visible in FPV, supporting safer teleoperation in unknown environments.
@@ II Related Work
Third-person limitations. Prior work found that third-person viewpoints can degrade navigation performance, reduce spatial orientation, and weaken embodiment in immersive tasks [25]. We consider a different use case, teleoperation of a physical drone. Here TPV complements egocentric cues by revealing lateral and rear hazards that FPV often hides. We therefore evaluate TPV against an FPV baseline and plan a larger study to identify when TPV helps and when it hurts.
@@ III-A Hexacopter Model
To localize the hexacopter motion, we define a global coordinate system that is fixed on the ground with orthonormal base vectors $\hat{\mathbf{e}}_{1}$, $\hat{\mathbf{e}}_{2}$, and $\hat{\mathbf{e}}_{3}$. We also define a local (body) coordinate system that is fixed on the hexacopter body, with orthonormal base vectors $\hat{\mathbf{c}}_{1}$, $\hat{\mathbf{c}}_{2}$, and $\hat{\mathbf{c}}_{3}$. We can use the 3213-2-1 Euler angle standard to characterize the orientation of the hexacopter at any time, where we use $\phi$, $\theta$, and $\psi$ to denote the hexacopter’s roll, pitch, and yaw angles, respectively. The angular velocity of the hexacopter is expressed with respect to the body coordinate system by
@@ III-B Controller
[…]
@@
[…]
@@ IV-A Hardware Setup
Figure 4: Second pilot conducting VR teleoperation with the Meta Quest 3 inside the safety enclosure (TPV/FPV trials).
@@ IV-B Software Setup
Latency measurement. At each frame, the VR server publishes a timestamped PoseStamped on /vrserver/tracking. The client echoes the most recent message back on /vrserver/tracking_echo. We compute round-trip time (RTT) between transmit and echo receipt and report a one-way estimate as RTT/2/2. Rendering on the headset is small relative to transport and treated as negligible. One-way statistics (median, mean, P95, max, and fraction >100>100 ms) are computed offline from the echo logs for pre/post comparisons.
@@
Figure 6: Comparison of latency with commercial offering.
Figure 7: One-way tracking latency over time estimated as RTT/2. The shaded interval marks a sustained mapping burst where variance rises. The two highest excursions occur at system startup and at a terminal emergency stop issued via the PX4 controller at the end of the flight. Operators can pause mapping or switch to FPV when tail latency grows (see Table IV).
Figure 8: Post-optimization one-way tracking latency over time (RTT/2). Dashed line shows 100 ms. Full run: median 16.1 ms, mean 29.8 ms, P95 86.0 ms, max 497.4 ms, and 3.16% of frames >100>100 ms. Steady state, excluding the first 10 s and frames >250>250 ms: median 15.9 ms, mean 24.2 ms, P95 74.6 ms, max 246.4 ms, and 1.47% >100>100 ms. The two largest spikes are attributable to system startup and a terminal PX4 emergency stop; neither reflects steady-state operation.
Figure 9: Post-optimization latency summary. Full run: median 16.1 ms, mean 29.8 ms, P95 86.0 ms, max 497.4 ms, and 3.16% >100>100 ms. Steady state, excluding first 10 s and frames >250>250 ms: median 15.9 ms, mean 24.2 ms, P95 74.6 ms, max 246.4 ms, and 1.47% >100>100 ms.
(a) Task time
(b) Minimum obstacle distance
Figure 10: FPV vs TPV performance in matched trials. Top, task time: FPV $55\pm 6$ s, TPV $57\pm 7$ s. Bottom, minimum distance: FPV 0.62 m, TPV 0.82 m.
Figure 11: Screenshot of partially reconstructed environment, hexacopter, and video stream.
@@
    Portability and flexibility, WebXR works on any compatible headset.
@@
    Bandwidth utilization, compressed video and a binary mesh format reduce load.
@@
    Security, transport is encrypted with TLS.
IV-C Robustness and Operational Fallbacks
Manual mapping throttle. When transient load increases, the operator can disable mapping using the built-in toggle (ROS2 service /vrserver/mapping, controller select). In this mode, only pose and video are streamed.
FPV-only fallback. If persistent congestion or mapping load impairs readability, the client can run in FPV mode by launching the page with ?mode=fpv. This shows a full-screen video feed while still rendering vehicle pose.
Figure 12: FPV camera stream rendered in the VR client. FPV-only fallback prioritizes the video feed while maintaining pose cues.
State indicators. The backend publishes tracking and mapping state on /vrserver/status; the front end surfaces status cues. Operator SOP. If one-way latency exceeds about 100 ms or state is non-OK for more than about 2 s, the sequence is: disable mapping, if needed switch to FPV, then re-enable mapping once stable.
IV-D Limitations
Our current evaluation is constrained to an indoor flight volume because the pose backbone relies on a Vicon motion-capture system. This infrastructure provides high-rate, high-accuracy ground truth but precludes outdoor trials, where GNSS availability, wind, lighting variability, and larger operational baselines may affect mapping quality and link behavior. The laboratory is a controlled facility with restricted access, which limits the participant pool. Two trained pilots flew the protocol. The sample size is small, so results should be interpreted as evidence of technical feasibility in controlled conditions rather than population-level usability. Future work will migrate to a VIO+GNSS/RTK outdoor stack and conduct an IRB-approved multi-participant study in open test ranges.
@@ V Experimental Results
Two trained pilots completed the protocol. We evaluated TPV (third-person VR with live mapping) and an FPV baseline (video-only). Each pilot executed: takeoff $\rightarrow$ approach a wall and stop at commanded standoff $\text{d}_{\mathrm{cmd}}] \rightarrow$ return $\rightarrow$ land. For quantitative analysis we report repeated trials from one pilot (N=8N{=}8) to control within-subject variability.
Metrics. We logged task time, minimum obstacle distance, contacts, path length, and velocity smoothness (acceleration $\ell_{2}$). Latency was measured via timestamped ROS2 pings echoed from the headset, yielding one-way estimates (median, mean, P95, max, and fraction >100>100 ms).
TPV vs FPV Performance. Table III summarizes performance. Task time was comparable (FPV $55{\pm}6$ s vs TPV $57{\pm}7$ s). Minimum obstacle distance improved in TPV (FPV $0.62$ m $\rightarrow$ TPV $0.82$ m, +0.20+0.20 m), and contacts were lower (FPV $1/4$ vs TPV $0/4$). Path length and smoothness differed by $5\%$.
TABLE III: TPV vs FPV (single pilot, repeated trials; reported metrics)
Condition Time (s) Min Dist (m) Contacts Smoothness
FPV $55{\pm}6$ 0.62 1/4 baseline
TPV $57{\pm}7$ 0.82 0/4 within 5%
V-A Latency after Optimizations
Using the timestamped echo method, the post-optimization one-way latency for the full run is: median 16.1 ms, mean 29.8 ms, P95 86.0 ms, max 497.4 ms, and 3.16% >100>100 ms. Because the largest excursions occur at startup and at a terminal PX4 emergency stop, we also report a steady-state slice excluding the first 10 s and frames >250>250 ms: median 15.9 ms, mean 24.2 ms, P95 74.6 ms, max 246.4 ms, and 1.47% >100>100 ms.
TABLE IV: One-way tracking latency (ms), before vs. after optimizations
Condition Median Mean P95 Max %>>100
Pre-opt (full) 28.7 41.9 105.0 507.0 6.0
Post-opt (full) 16.1 29.8 86.0 497.4 3.16
Post-opt (steady-state) 15.9 24.2 74.6 246.4 1.47
Context vs Commercial Latency As in Fig. 6, the post-optimization median remains below commonly cited real-time perception thresholds (Fig. 9) and comparable to commercial FPV offerings, while P95 and tail behavior improve when mapping bursts are avoided.